OpenStudy (anonymous):

What is the equation of the line that passes through ( 1,5) and (4,11)?

OpenStudy (anonymous):

first find the slope then use the point-slope formula

OpenStudy (anonymous):

the slope is y2-y1/x2-x1 = 11-5/4-1 = 6/3 =2

OpenStudy (anonymous):

The point-slope formula is y-y1 = m(x-x1) so pick one of the two points and plug in for x1 and yi and use the slope,2, for m and that is your equation

OpenStudy (anonymous):

My equation is y=2x+3.. is this correct? @asdfvg

OpenStudy (anonymous):

yup :)